OpenStudy (anonymous):

factor completely 80c^2+200cf+125f^2

OpenStudy (mertsj):

Factor out 5: 5(16c^2 + 40cf + 25f^2) = 5(4c+5f)^2

OpenStudy (anonymous):

huh

OpenStudy (mertsj):

Do you understand about factoring out the 5?

OpenStudy (anonymous):

a little but not a lot

OpenStudy (mertsj):

5 is a factor of 80, 200, and 125. Therefore it must be factored out using the distributive property: \[80c ^{2}+200cf +125f ^{2} =5(16c ^{2}+40cf + 25f ^{2}\]

OpenStudy (mertsj):

Do you understand that?

OpenStudy (anonymous):

yeah

OpenStudy (mertsj):

Ok. Then consider \[16c ^{2}+ 40cf + 25f ^{2} \] That is a trinomial and came from multiplying two binomials. It is our job to find the two binomials. To get 16c^2 we could multiply 4c by 4c. Let's try that. (4c )(4c ) To get 25f^2 on the end we could have 5f times 5f. Let's try that and see if we get the correct middle term of 40cf (4c+5f)(4c+5f). The outer is 20cf and the inner is 20 cf. If we add those, we get 40 cf. So that seems to work and the factored form is therefore 5(4c+5f)(4c+5f) of 5(4c+5f)^2

OpenStudy (anonymous):

so then i would be at the answer of 5(4c+5f)^2 right now where do i go from there

OpenStudy (mertsj):

That's it. It's completely factored.

OpenStudy (anonymous):

so in my problem the answer was sort of there in a way right with the factoring of 5 right

OpenStudy (mertsj):

That was the first step. That is always the first step in any factoring problem. To look for a common factor. Sometimes there will be a common factor but not always. If there is a common factor, the first step is to factor it out using the distributive property.

OpenStudy (mertsj):

Then the second step is to examine what remains after factoring out the common factor. Sometimes it will be factorable but not always.

OpenStudy (anonymous):

you explained that really easy for me in a way,thank you